A.) What is the pressure drop (in N/m2) due to the Bernoulli effect as water goes into a 3.10 cm diameter nozzle from a 8.80 cm diameter fire hose while carrying a flow of 36.0 L/s?

B.) To what maximum height (in m) above the nozzle can this water rise? (The actual height will be significantly smaller due to air resistance.)
Answer #1

here,

diameter of nozzle , d1 = 3.1 cm = 0.031 m

diameter of pipe , d2 = 8.8 cm = 0.088 m

flow speed , V = 36 L/s = 0.036 m^3 /s

the speed of water in nozzle ,v1 = V /(pi * (d1/2)^2)

v1 = 0.036 /( pi * (0.031/2)^2)

v1 = 47.72 m/s

the speed of water in nozzle , v2 = V /(pi * (d2/2)^2)

v2 = 0.036 /( pi * (0.088/2)^2)

v2 = 5.92 m/s

the pressure drop , dP = 0.5 * density of water * ( v1^2 - v2^2)

dP = 0.5 * 1000 * ( 47.72^2 - 5.92^2)

dP = 1.12 * 10^6 Pa

B)

the maximum height above the nozzle can this water rise , hm = v2^2 /(2g)

hm = 5.92^2 /(2*9.81) m

hm = 1.79 m